# Graph Coloring for Multi-Task Learning

### v2

@@ 7.1 Overall Performance Improvements
Overall, the conflict-aware approaches improve over the uniform baseline by 10%-20% on CIFAR-10 and by 7% on MM-IMDb. This reinforces the idea that grouping tasks according to measured interference is more effective than treating all tasks equally at every update. On NYUv2, we see similar improvements across all the metrics. These results suggest that the scheduler’s graph coloring cleanly separates high-conflict tasks, preserving the projection or LR-balancing advantages (stemming from PCGrad’s gradient projection and AdaTask’s learning‑rate adaptation, respectively) while removing residual interference.
@@ 7.2 Ablation Study on Scheduler Design
Our ablation study (Section 6.4) further highlights the importance of how SON-GOKU is designed. The results show that: (1) Dynamic recoloring matters. Static One-Shot underperforms the full scheduler on most metrics, indicating that task relations change enough during training that frozen groups become stale. This supports the need for periodic refresh; and (2) History smoothing is very important. Single-Step is markedly worse across datasets, consistent with our claim that per-batch cosines are too noisy to define stable groups. This aligns with our recovery analysis, which requires concentration of EMA gradients within a refresh window.
7.2.1 Interpretation of Ablation Study Results
Dynamic recoloring lets the schedule track how task interactions change over time, so partitions do not go stale. Averaging gradients over a short history makes the conflict signal less noisy, which results in more stable groups and reliable progress (Section 5.1 and 5.4). Together, these choices satisfy the conditions we use in our theory. They create low-conflict groups that ensure safe per-step descent (Section 5.15.2) and provide enough concentration to recover the population conflict graph within each refresh (Section 5.4).
@@ 7.3.1 Optimizer-Task Alignment
We believe that this stems from unique differences in the features of classification and regression-based models. For example, cross-entropy gradients near decision boundaries tend to be bursty and high in variance (Shrivastava et al., 2016; Lin et al., 2017; Hoffer et al., 2017). By scaling each task’s step size according to its running gradient norm, AdaTask smooths out these spikes.
@@ 7.3.2 Synergy Between Scheduling and Baselines
Finally, by computing interference over a window, the scheduler smooths out gradient fluctuations. This prevents the erratic schedule changes that projection-only grouping methods have been shown to face (Yu et al., 2020; Shi et al., 2023; Zhang et al., 2024), thereby better stabilizing convergence.
@@ 7.4 Speed and Tradeoffs
The proposed scheduler has a time complexity of $\Theta\!\big(K^{2}d/R\big)$ amortized per training step (Section 4.5). Table 2 shows near-linear growth over this range of $K$ at R=32R{=}32, reflecting sparsity in the graphs and batched cosine computation. SON-GOKU’s time rises from around 2 seconds ($K=3$) to 12 seconds ($K=40$), remaining far below methods that perform heavy per-step conflict handling. For example, PCGrad, FairGrad, and Nash-MTL increase steeply with $K$. In contrast, FAMO and AdaTask are among the fastest and largely flat with $K$, as expected from their constant overhead.
These contrasts demonstrate the tradeoffs between speed and fidelity to task interference. Faster methods like FAMO minimize overhead, while methods that model conflicts can improve accuracy. These tradeoffs have to be assessed on a case-by-case basis, based on values that factor into each approach’s time complexity and the importance of training speed versus performance on an application.
@@ B.2 Assumptions
We adopt the following assumptions, which are standard in analyses of stochastic-gradient methods and verifiable in practice (see, e.g., Robbins & Monro 1951; Kushner & Yin 2003; Nemirovski et al. 2009; Bottou et al. 2018; Wainwright 2019; for concentration of geometrically weighted and mixing sequences, see Merlevède et al. 2011; De la Pena et al. 2009).
@@ C.4 Scheduled and aggregated gap at a common linearization
Using $\|H(\cdot)\|_{\mathrm{op}}\leq L$ and Cauchy-Schwarz (Steele, 2004)
@@ C.7.1 $L$-smoothness
This is the same regularity used throughout the main paper and in our baselines. Each task loss we optimize is $L_{i}$-smooth, so the overall objective is $L$-smooth. We only use this to apply the standard smoothness (descent) inequality (Nesterov, 2004; Beck, 2017).
@@ C.7.4 PL and strong convexity
We invoke PL only to convert a per-refresh decrease into a standard contraction factor. We do not require global strong convexity. A local PL inequality around the iterates is enough, which is commonly observed after warm-up and annealing we already use (Karimi et al., 2016; Zhou et al., 2021; Liu, 2025).
@@ D.1 Notation
D.2 Per-refresh complexity (time and space)
@@
denote the matrix whose $i$-th row $m_{i}^{\top}$ is the EMA for task $i$. A refresh first updates these rows through a scalar EMA rule
@@
using the most recent probe (or reused) gradient $g_{i}$. It then constructs the cosine-similarity matrix
$\widehat{C}=\widetilde{M}\widetilde{M}^{\top}$ (66)
where $\widetilde{M}$ is the row-normalized version of $M$. It thresholds $\widehat{C}$ at $\tau$ to obtain the conflict adjacency. Finally, it applies a graph-coloring routine to the resulting simple graph (Welsh & Powell, 1967).
@@
The construction of $\widehat{C}$ proceeds by normalizing each row of $M$ and then multiplying $\widetilde{M}$ by its transpose. Row normalization touches each entry exactly once and therefore costs $\Theta(Kd)$ time. The Gram product $\widetilde{M}\widetilde{M}^{\top}$ consists of $K^{2}$ dot products of length $d$, which is $\Theta(K^{2}d)$ time (Kågström et al., 1998). The cosine matrix itself occupies $K^{2}$ entries. If it is retained after thresholding, it uses $\Theta(K^{2})$ space. If dropped right after graph construction, that $\Theta(K^{2})$ storage is only temporary.
@@
Proposition 5 (Per-refresh scheduler overhead).
Under the standard RAM model with dense matrix multiplication costed as $\Theta(K^{2}d)$, the time required by a single scheduler refresh is
$T_{\text{refresh}}\;=\;\Theta(Kd)\;+\;\Theta(K^{2}d)\;+\;\Theta(K^{2})\;+\;O(K^{2})\;=\;\Theta(K^{2}d),$ (67)
@@
$S_{\text{refresh}}\;=\;\Theta(Kd)\;+\;\Theta(K^{2}),$ (68)
where the $\Theta(K^{2})$ term is transient if $\widehat{C}$ is not retained after coloring.
@@
The EMA update costs $\Theta(Kd)$ by a direct count of coordinate-wise multiply-adds. Row normalization also costs $\Theta(Kd)$. The Gram matrix requires $K^{2}$ inner products of length $d$, which is $\Theta(K^{2}d)$. This term dominates $\Theta(Kd)$. Thresholding scans $O(K^{2})$ entries and is therefore $\Theta(K^{2})$. The greedy coloring performs a sort of $K$ keys and then assigns at most one color per edge incident on the current vertex, which is $O(K^{2})$ in the worst case. This is dominated by $\Theta(K^{2}d)$ whenever $d\geq 1$. Summing these contributions and absorbing lower-order terms yields $T_{\text{refresh}}=\Theta(K^{2}d)$. The EMA matrix occupies $\Theta(Kd)$ memory, and storing $\widehat{C}$ uses $\Theta(K^{2})$. But if $\widehat{C}$ is discarded immediately after thresholding, only $\Theta(Kd)$ remains. ∎
@@ D.3 Amortized cost over training
$\frac{1}{R}\,T_{\text{refresh}}\;=\;\ \frac{1}{R}\,\Theta(K^{2}d)\$(70)
If probe gradients are computed only at refreshes, their contribution $K\mathsf{G}$ per refresh adds $\frac{1}{R}\Theta(K\mathsf{G})$ to the amortized time per step. If, instead, the training loop already computes task-wise gradients each step and these are reused to update the EMAs, then the probe term is absent and the amortized scheduler overhead remains $\frac{1}{R}\Theta(K^{2}d)$.
@@ D.4 Conditions for negligible overhead
$C_{\text{sched}}=\frac{a}{R}\,K^{2}d\qquad\text{and}\qquad C_{\text{probe}}=\frac{b}{R}\,K\,\mathsf{G},$ (71)
@@
$\frac{C_{\text{sched}}}{C_{\text{probe}}}=\frac{a}{b}\,\frac{K^{2}d}{K\,\mathsf{G}}=\frac{a}{b}\,\frac{Kd}{\mathsf{G}}.$ (72)
@@
$\frac{C_{\text{sched}}}{C_{\text{probe}}}\leq\varepsilon\quad\text{for some }0<\varepsilon\ll 1\quad\Longleftrightarrow\quad Kd\;\leq\;\frac{b}{a}\,\varepsilon\,\mathsf{G}.$ (73)
@@ D.5.1 Random projections
We replace the EMA matrix $M\in\mathbb{R}^{K\times d}$ by a lower-dimensional sketch $\widetilde{M}=MR$ with $R\in\mathbb{R}^{d\times r}$ and $r\ll d$ (Dasgupta & Gupta, 2003). The sketching multiply costs $O(Kdr)$ and the cosine Gram becomes $O(K^{2}r)$ instead of $\Theta(K^{2}d)$. Storage for the sketched EMAs is $O(Kr)$. By the Johnson-Lindenstrauss (JL) random projection guarantee, if we map the $K$ task-EMA vectors from $\mathbb{R}^{d}$ to $\mathbb{R}^{r}$ using a suitable random matrix with $r=\Theta(\epsilon^{-2}\log K)$, then after row normalization all pairwise inner products (hence cosines) are preserved within $\pm\epsilon$ with high probability. We assume a uniform row-norm floor $\min_{i}\|m_{i}\|\geq m_{0}>0$ (which can be enforced in practice by skipping tasks with $\|m_{i}\|<\nu\ll m_{0}$) so cosine errors remain controlled. Choosing $\epsilon<\gamma$, where $\gamma$ is the cosine margin from the recovery analysis, ensures that every pair remains on the same side of the threshold $\tau$. Therefore the set $\{(i,j):\widehat{C}_{ij}<-\tau\}$ and the resulting coloring are unchanged with high probability.

### v4

@@ 7.1 Overall Performance Improvements
Overall, the conflict-aware approaches improve over the uniform baseline by 10%-20% on CIFAR-10 and by 7% on MM-IMDb, indicating that grouping tasks according to measured interference is more effective than treating all tasks equally at every update. On NYUv2, we see similar improvements across all the metrics. These results suggest that the scheduler’s graph coloring cleanly separates high-conflict tasks, preserving the projection or LR-balancing advantages (stemming from PCGrad’s gradient projection and AdaTask’s learning‑rate adaptation, respectively) while removing residual interference (see App. S for grouping patterns at training time and more analyses). As we evaluated across diverse tasks and datasets, our results also demonstrate clear improvements in generalization.
@@ 7.2 Ablation Study on Scheduler Design
We evaluate nine controlled ablations of six types: (i) Static One-Shot Coloring, which runs greedy graph coloring once at the start of training and then freezes the groups, testing dependence on dynamic recoloring as gradients change; (ii) Single-Step Conflict Estimation, which sets the history length to $H=1$ so each recoloring uses only the most recent mini batch, testing the importance of averaging conflict statistics over time; (iii) Threshold Graph (baseline), which connects tasks $i$ and $j$ when the smoothed cosine $\hat{s}_{ij}(t)$ falls below a global threshold $\tau(t)$; (iv) kNN-Symmetric Graph, which connects each task to its $m$ most conflicting neighbors and then symmetrizes the edges, enforcing roughly fixed degree per task and comparing local degree control against the global threshold rule; (v) Signed-Only Graph, which adds an edge only if $\hat{s}_{ij}(t)<0$, yielding a very sparse graph and ignoring moderate (but potentially harmful) conflicts; and (vi) Quantile Threshold Graph, which at each refresh sets $\tau(t)$ so that only the worst $p\%$ of cosine values are treated as conflicting, keeping edge density approximately stable and testing an adaptive cutoff versus a fixed global threshold. We evaluate each graph rule under two settings. In the fixed $\tau$ setting, all rules share the same $\tau(t)$ schedule used in the main experiments. In the density-matched setting, we adjust the hyperparameters of each rule so that all graphs have approximately the same edge density at each refresh. This isolates the effect of which pairs are marked as conflicting, rather than how many edges are present. We go into much further detail regarding the ablation in Appendix K.3.
These ablations directly test the assumptions behind SON-GOKU. Static One-Shot, which freezes groups, consistently underperforms the full scheduler on most metrics, indicating that task relations change enough during training that dynamic recoloring is needed to maintain $\tau$-compatibility as gradients drift (Sections 5.15.2). Single-Step, which uses $H=1$, is clearly worse across datasets, matching our claim that batch cosines are too noisy. Instead, averaging conflict statistics over short history windows provides the clean information needed for accurate graph recovery (Section 5.4). Among graph constructions, simple threshold and quantile rules (and their density-matched variants) perform similarly well, suggesting that any approach that reliably isolates the worst conflicting pairs is sufficient. In contrast, Signed-Only and kNN-Symmetric, which ignore conflict magnitude or have purely local degree control, degrade performance more noticeably, especially on NYUv2 and the tabular benchmarks. Overall, the best performing configurations are precisely those that match the descent and recovery conditions analyzed in Sections 5.15.2 and 5.4.
@@ 7.3.1 Optimizer-Task Alignment
We believe that this stems from unique differences in the features of classification and regression-based models. For example, cross-entropy gradients near decision boundaries tend to be bursty and high in variance (Shrivastava2016OHEM; Lin2017FocalLoss; Hoffer2017TrainLonger). By scaling each task’s step size according to its running gradient norm, AdaTask smooths out these spikes.
@@ 7.3.2 Synergy Between Scheduling and Baselines
Finally, by computing interference over a window, the scheduler smooths out gradient fluctuations. This prevents the erratic schedule changes that projection-only grouping methods have been shown to face (Yu2020PCGrad; Shi2023Recon; Zhang2024Proactive), thereby better stabilizing convergence.
7.3.3 SON-GOKU’s Ability to Create Generalizable Models
While our guarantees in Section 5 and Appendices BF are stated in optimization terms, they also directly increase gradient coherence and limit destructive interference in ways that are known to favor generalization to unseen data. Section 5.1 shows that the aggregated group gradient remains aligned with descent and that intra-group gradient conflict is explicitly limited by $\tau$ and $|S_{t}|$. Section 5.3 then compares two ways to apply the same gradients during a refresh, either a single mixed update or a scheduled sequence of group updates. Together, these analyses imply that each step in SON-GOKU provides more informative signals and less interference, or, equivalently, a higher gradient-to-noise ratio (sun2023unleashing; Fan2023MaxGNR; mccandlish2018empirical). Building on this, Section 5.4 shows that SON-GOKU’s estimated conflict graph recovers the population structure with high probability, so the schedule repeatedly updates clusters of related tasks rather than conflicting tasks. By enforcing positive affinity within groups, SON-GOKU is able to train related tasks together. This enables effective sharing of model parameters across different tasks, reducing the complexity of the model and increasing sample efficiency (Caruana1997; Argyriou2007MultiTaskFeature; 9336293). With this alongside a high gradient-to-noise signal ratio, SON-GOKU theoretically generalizes across many different datasets, domains, and distributions and can perform well even under non-ideal conditions (e.g., noisy labels, class or task imbalance, distribution shift, etc.) (michalkiewicz2023domain). Our ablation results (Table 1) demonstrate that variants without dynamic recoloring or history averaging perform worse, indicating accurate and low-conflict grouping is essential.
@@ 7.4 Speed and Tradeoffs
SON-GOKU has a time complexity of $O\!\big(Kr(d+K)/R\big)$ (Section 4.5) amortized per training step (Section 4.5). Table 2 shows near-linear growth over this range of $K$ at R=32R{=}32, reflecting sparsity in the graphs and batched cosine computation. SON-GOKU’s time rises from around 2 seconds ($K=3$) to 12 seconds ($K=40$), remaining far below methods that perform heavy conflict handling. For example, PCGrad, FairGrad, and Nash-MTL increase steeply with $K$. In contrast, FAMO and AdaTask are among the fastest and largely flat with $K$, as expected from their constant overhead.
SON-GOKU is also memory efficient, with an only incremental memory footprint that scales with the number of tasks $K$, not the parameter dimension $d$. The scheduler’s peak memory during a refresh step is $O(K^{2}{+}Kr)$ and the persistent state between refreshes is $O(K^{2})$ (see Appendix N for further theoretical and experimental analysis). By contrast, methods that retain $K$ full gradients require $O(Kd)$ additional memory. This implies that, on larger backbones (high $d$), SON-GOKU’s memory overhead is modest and grows mainly with the task count $K$, rather than with model size.
These contrasts demonstrate the tradeoffs between speed and fidelity to task interference. Faster methods like FAMO minimize overhead, while methods that model conflicts can improve accuracy. These tradeoffs have to be assessed on a case-by-case basis, based on values that factor into each approach’s time complexity and the importance of training speed versus performance.
@@ 8 Conclusion
9 Reproducibility Statement
We provide a clean code repository for reproducibility in the supplementary materials, and this is also provided in an online (de-identified) Git repository. The scripts in this repository contain functionality for downloading, loading, and preprocessing all datasets used in training. The code also includes implementations for SON-GOKU and all 10 of its ablations. We provide clear and easy-to-use training scripts with pre-configured parameters, allowing for reproduction of the exact experiments used across all datasets. We provide further details regarding empirical experiments and evaluation in Appendices JL. To make our theoretical analysis easier to follow and more transparent, we provide highly detailed descriptions of assumptions, propositions, and proofs that could not fit in the main text in Appendices BI. Furthermore, to make effective real-world deployment easier, we provide practical guidance regarding SON-GOKU in Appendices D, N, and O.
@@ B.2 Assumptions
We adopt the following assumptions, which are standard in analyses of stochastic-gradient methods and verifiable in practice (see, e.g., robbins1951stochastic; kushner2003sa; Nemirovski2009; bottou2018optimization; wainwright2019hds; for concentration of geometrically weighted and mixing sequences, see merlevede2011bernstein; delapena2009self).
@@ C.4 Scheduled and aggregated gap at a common linearization
Using $\|H(\cdot)\|_{\mathrm{op}}\leq L$ and Cauchy-Schwarz (steele2004cauchy)
@@ C.7.1 $L$-smoothness
This is the same regularity used throughout the main paper and in our baselines. Each task loss we optimize is $L_{i}$-smooth, so the overall objective is $L$-smooth. We only use this to apply the standard smoothness (descent) inequality (Nesterov2004; Beck2017).
@@ C.7.4 PL and strong convexity
We invoke PL only to convert a per-refresh decrease into a standard contraction factor. We do not require global strong convexity. A local PL inequality around the iterates is enough, which is commonly observed after warm-up and annealing we already use (Karimi2016PL; Zhou2021Local; Liu2025Warmup).
@@ D.1 Notation
  • $r\in\mathbb{N}$ – dimension of the sketch space used for cosine computation (number of columns of the random projection matrix)
D.2 Per-refresh complexity
@@
denote the matrix whose $i$-th row $m_{i}^{\top}$ is the EMA for task $i$. We fix a random matrix $R\in\mathbb{R}^{d\times r}$ with $r\ll d$ and form a lower dimensional sketch
$M_{f}:=MR\in\mathbb{R}^{K\times r}.$ (65)
A refresh first updates these rows through a scalar EMA rule
@@
using the most recent probe (or reused) gradient $g_{i}$. It then constructs the cosine-similarity matrix in the sketch space
$\widehat{C}=\widetilde{M}_{f}\widetilde{M}_{f}^{\top}$ (67)
where $\widetilde{M}_{f}$ is the row-normalized version of $M_{f}$. It thresholds $\widehat{C}$ at $\tau$ to obtain the conflict adjacency. Finally, it applies a graph-coloring routine to the resulting simple graph (Welsh1967).
@@
The construction of $\widehat{C}$ in the sketched space proceeds in three stages: (i) forming the sketch $M_{f}=MR$, (ii) normalizing each row of $M_{f}$, and (iii) multiplying $\widetilde{M}_{f}$ by its transpose. The sketching multiply touches every entry of $M$ and $R$ and therefore costs $O(Kdr)$ time. Row normalization touches each entry of $M_{f}$ exactly once and therefore costs $\Theta(Kr)$ time. The Gram product $\widetilde{M}_{f}\widetilde{M}_{f}^{\top}$ consists of $K^{2}$ dot products of length $r$, which is $O(K^{2}r)$ time (kaagstrom1998gemm). The cosine matrix itself occupies $K^{2}$ entries. If it is retained after thresholding, it uses $\Theta(K^{2})$ space. If dropped right after graph construction, that $\Theta(K^{2})$ storage is only temporary.
@@
Proposition 5 (Per-refresh scheduler overhead with random projections).
Under the standard RAM model with dense matrix multiplication in the sketch space costed as $O(K^{2}r)$, the time required by a single scheduler refresh is
$T_{\text{refresh}}\;=\;\Theta(Kd)\;+\;O(Kdr)\;+\;O(K^{2}r)\;+\;O(K^{2})\;=\;O(Kdr+K^{2}r),$ (68)
@@
$S_{\text{refresh}}=\Theta(Kd)+\Theta(Kr)+\Theta(K^{2}),$ (69)
where the $\Theta(K^{2})$ term is transient if $C$ is not retained after coloring and the $\Theta(Kr)$ term is transient if the sketch $M_{f}$ is discarded between refreshes and recomputed from $M$.
@@
The EMA update costs $\Theta(Kd)$ by a direct count of coordinate-wise multiplication and addition. Forming the random projection sketch $M_{f}=MR$ touches each entry of $M$ and $R$, and therefore costs $O(Kdr)$. Row-normalizing $M_{f}$ then costs $\Theta(Kr)$, since it processes all $Kr$ entries once.
In the sketched space, the Gram matrix $\widehat{C}=\widetilde{M}_{f}\widetilde{M}_{f}^{\top}$ requires $K^{2}$ inner products of length $r$, which is $O(K^{2}r)$ time. Thresholding scans $O(K^{2})$ entries and is therefore $\Theta(K^{2})$. The greedy coloring performs a sort of $K$ keys and then assigns at most one color per edge incident on the current vertex, which is $O(K^{2})$ in the worst case. These $\Theta(K^{2})$ terms are dominated by the $O(Kdr)$ and $O(K^{2}r)$ contributions once $r\geq 1$ and $K$ is nontrivial.
Summing these contributions and absorbing lower-order terms yields $T_{\text{refresh}}\;=\;\Theta(Kd)+O(Kdr)+O(K^{2}r)+O(K^{2})\;=\;O(Kdr+K^{2}r).$
@@ D.3 Amortized cost over training
$\frac{1}{R}T_{\text{refresh}}=\frac{1}{R}\,O(Kdr+K^{2}r)=O\!\left(\frac{Kdr+K^{2}r}{R}\right).$ (71)
If probe gradients are computed only at refreshes, their contribution $K\mathsf{G}$ per refresh adds $\frac{1}{R}\Theta(K\mathsf{G})$ to the amortized time per step. If, instead, the training loop already computes task-wise gradients each step and these are reused to update the EMAs, then the probe term is absent and the amortized scheduler overhead remains $O\!\big((Kdr+K^{2}r)/R\big)$.
@@ D.4 Conditions for negligible overhead
$C_{\text{sched}}=\frac{a}{R}\,(Kdr+K^{2}r),\qquad C_{\text{probe}}=\frac{b}{R}\,KG,$ (72)
@@
$\frac{C_{\text{sched}}}{C_{\text{probe}}}=\frac{a}{b}\,\frac{Kdr+K^{2}r}{KG}=\frac{ar}{b}\,\frac{d+K}{G}.$ (73)
@@
$\frac{C_{\text{sched}}}{C_{\text{probe}}}\leq\varepsilon\quad\Longleftrightarrow\quad r\,(d+K)\;\leq\;\frac{b}{a}\,\varepsilon\,G.$ (74)
@@ D.5.1 Random projections
We replace the EMA matrix $M\in\mathbb{R}^{K\times d}$ by a lower-dimensional sketch $\widetilde{M}=MR$ with $R\in\mathbb{R}^{d\times r}$ and $r\ll d$ (DasguptaGupta2003). The sketching multiply costs $O(Kdr)$ and the cosine Gram becomes $O(K^{2}r)$ instead of $\Theta(K^{2}d)$. Storage for the sketched EMAs is $O(Kr)$. By the Johnson-Lindenstrauss (JL) random projection guarantee, if we map the $K$ task-EMA vectors from $\mathbb{R}^{d}$ to $\mathbb{R}^{r}$ using a suitable random matrix with $r=\Theta(\epsilon^{-2}\log K)$, then after row normalization all pairwise inner products (hence cosines) are preserved within $\pm\epsilon$ with high probability. We assume a uniform row-norm floor $\min_{i}\|m_{i}\|\geq m_{0}>0$ (which can be enforced in practice by skipping tasks with $\|m_{i}\|<\nu\ll m_{0}$) so cosine errors remain controlled. Choosing $\epsilon<\gamma$, where $\gamma$ is the cosine margin from the recovery analysis, ensures that every pair remains on the same side of the threshold $\tau$. Therefore the set $\{(i,j):\widehat{C}_{ij}<-\tau\}$ and the resulting coloring are unchanged with high probability.